When 𝐼 = 10, π‘Ž = 2 and 𝑛 = 5, calculate the value of 𝑑 in the formula below:

𝐼 = π‘Ž + (𝑛 βˆ’ 1) 𝑑

[1] 𝑑 = 2

[2] 𝑑 = 3

[3] 𝑑 =48

[4] 𝑑 = 8


1
Expert's answer
2022-04-12T11:30:57-0400

𝐼 = π‘Ž + (𝑛 βˆ’ 1) 𝑑

d = (I - a) / (n -1) = (10-2) / (5 -1) = 2


[1] 𝑑 = 2